OpenStudy (anonymous):

Could someone walk me through this question and explain to me what happens. I have been working on it for 7 hours. I don't understand my prof and I might fail the class and I want to cry. Its solving a quadratic equation that had complex numbers

OpenStudy (anonymous):

\[2x^2-ix-2+i=0\]

OpenStudy (anonymous):

:/ Sorry... I can't help..

OpenStudy (anonymous):

:'(

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

tried using the quadratic formula yet?

OpenStudy (anonymous):

Phew! GO JDOE!

OpenStudy (anonymous):

I tried it, i solved it but i got the wrong answers. I tried it over and over again. I looked up videos, still wrong answer

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} {\color{red}{ 2}}x^2&{\color{blue}{ -i}}x&{\color{green}{ -2+i}}=0\\ \uparrow &\ \uparrow &\quad \uparrow \\ {\color{red}{ a}}&\ {\color{blue}{ b}}&\quad {\color{green}{ c}} \end{array}\\ \quad \\ \bf x=\cfrac{i\pm\sqrt{i^2-4(2)(-2+i)}}{2(2)}\implies x=\cfrac{i\pm\sqrt{-1-8(-2+i)}}{2(2)}\\ \quad \\ x=\cfrac{i\pm\sqrt{-1+16-8i}}{2(2)}\implies x=\cfrac{i\pm\sqrt{15-8i}}{4}\)

OpenStudy (anonymous):

ok..got that

OpenStudy (jdoe0001):

well.. that's.... as far as I can see it going... what does your answer say?

OpenStudy (anonymous):

x=1 and x=-1+0.5i

OpenStudy (jdoe0001):

hmmm I don't see it simplifying ... to that

OpenStudy (anonymous):

:(

OpenStudy (anonymous):